OpenStudy (anonymous):

The time required to finish a test in normally distributed with a mean of 80 minutes and a standard deviation of 15 minutes. What is the probability that a student chosen at random will finish the test in more than 110 minutes?

OpenStudy (anonymous):

34% 16% 82% 2%

OpenStudy (anonymous):

\[P(X>110)=P\left(\frac{X-80}{15}>\frac{110-80}{15}\right)=P(Z>2)\]

OpenStudy (anonymous):

If you're not familiar with \(z\) tables and the computation of these sorts of probabilities, the other way to determine the value of \(P(Z>2)\) is to ask - what proportion of a population lies 2 standard deviations away from the mean?